contract the expressions. that is, use the properties of logarithms to write each expression coefficient of 1.
Ln 3 - 2 ln(6+9)
how do i work out this problem?

Answer by DrBeeee(684)

Given: (1) ln(3) - 2*ln(6+9) or
(2) ln(3) - 2*ln(15) or
using 2*ln(x) = ln(x)^2 we get
(3) ln(3) - ln(15)^2 or
(4) ln(3) - ln(225) or
using ln(a) - ln(b) = ln(a/b) we get
(5) ln(3/225) or
(6) ln(1/75) or
(7) ln(1) - ln(75) or
(8) 0 - ln(75) or
(9) -ln(75)
Answer: (1) evaluates to -ln(75) = -4.317...